### v2

@@ 5 Approximate-MMS with Copies: Additive Valuations
It has been observed, for the original MMS problem (without copies), that it is without loss of generality to consider ordered instances [16, 14]. Notably, where copies are allowed, the same transformation as in [16, 14] may no longer yield a valid allocation, see example in Appendix F. To address this, we impose an additional constraint on the allocation in the ordered instance (see Definition 5), and show that under this constraint, it suffices to consider ordered instances. We then devise algorithms that produce allocations adhering to this constraint.
In Section 5.2, we review some of the known reduction rules and introduce new ones. Then, in Section 5.3, we present a simple algorithm $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ which combines the idea of bag-filling and round robin to achieve an $\alpha$-MMS allocation with copies (see Algorithm 4). This algorithm is the building block for achieving $6/7$-MMS and $4/5$-MMS with copies in the sections that follow. In particular, in Sections 5.4 and 5.5, we obtain $6/7$-MMS with $\lfloor n/2\rfloor$ distinct copies and $4/5$-MMS with $\lfloor n/3\rfloor$ distinct copies, respectively, by first reducing the initial instance using appropriate valid reductions, and then applying $\mathsf{BagFill\mbox{-}RoundRobin}$.
@@ 5.1 Ordered Instance
Given an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, for every agent $i\in N$, let $\sigma_{i}:[m]\to[m]$ be a permutation that orders the goods according to $i$’s preferences, namely $\sigma_{i}(j)$ is the $j$th most valued good (breaking ties by index111any consistent tie breaking rule will do., and thus $v_{i}(\sigma_{i}(1))\geq v_{i}(\sigma_{i}(2))\geq\dots\geq v_{i}(\sigma_{i}(m))$.
As done by Bouveret and Lemaître [16], we construct an ordered instance $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ with a new set of “virtual goods”, $M^{\prime}=\{1,\dots,m\}$, and modified valuations, $\{v^{\prime}_{i}\}_{i\in N}$, such that $v^{\prime}_{i}(j)=v_{i}(\sigma_{i}(j))$ for any $j\in M^{\prime}$.
Definition 4 (Ordered Counterpart ([16])).
Given a fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, its ordered counterpart, $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$, is an instance with the same set of agents and a set of “virtual goods”, $M^{\prime}$, of the same cardinality as $M$. For each agent $i\in[n]$, and any $j\in M^{\prime}$, it holds that $v^{\prime}_{i}(j)=v_{i}(\sigma_{i}(j))$.
Observe that by definition, all agent in $\mathcal{I}^{\prime}$ have the same ordinal preferences, i.e., for every $i\in N$, $v^{\prime}_{i}(1)\geq v^{\prime}_{i}(2)\geq\dots\geq v^{\prime}_{i}(m)$. It was shown in [16] that to guarantee agent $i$ a value of at least $\alpha_{i}$ in an instance $\mathcal{I}$, it suffices to find an allocation providing this guarantee in its ordered counterpart $\mathcal{I}^{\prime}$. This is cast in the following proposition.
Proposition 2 (Reduction to ordered without copies [16]).
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance and let $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ be its ordered counterpart. Any allocation (without copies) $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ for $\mathcal{I}^{\prime}$ can be converted in poly-time to an allocation (without copies) $A=(A_{1},\dots,A_{n})$ for $\mathcal{I}$, such that for every agent $i$, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$.
@@
Since by iteration $j$ exactly $j-1$ items have been chosen, agent $i$ always chooses a good $g\in M$ such that $v_{i}(g)\geq v_{i}(\sigma_{i}(j))=v^{\prime}_{i}(j)$ and thus, by additivity, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$. ∎
We will show a similar result to Proposition 2, for allocations with $t$ distinct copies in which the top $t$ goods (and their copies) are allocated to $2t$ different agents.
Definition 5 (Simple Allocation).
@@
  • There is at most one copy of each good.
@@
Recall that agent $i$ cannot get two instances of the same good in $A^{\prime}_{i}$. [^2] Otherwise, and if additive valuations are extended so that two instances of the same good yield twice its value, then the same argument as in Proposition 2 would have implied the correctness of Proposition 3.
Proposition 3 (Reduction to ordered for simple allocations).
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance and let $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ be its ordered counterpart. Let $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ be a simple allocation for $\mathcal{I}^{\prime}$ with $t$ distinct copies. Then, $A^{\prime}$ can be converted in poly-time to an allocation for $\mathcal{I}$ with $t$ distinct copies, $A=(A_{1},\dots,A_{n})$, such that for every agent $i$, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$.
In order to prove Proposition 3, we presented the following useful lemma.
Lemma 7.
@@
If there exists a set $G\subset M\setminus X$ of $\ell$ distinct goods such that $v_{i}(X\cup\{g\})\geq\tau$ for all $g\in G$, then there exists a set $H\subset M\setminus Y$ of $\ell$ distinct goods such that $v_{i}(Y\cup\{h\})\geq\tau$ for all $h\in H$.
@@
For every $j\in[k]$, we add the (black) edge $(\mathcal{Y}_{j},\mathcal{X}_{j})$. In addition, for any good $g\in X\cap Y$, we add the (red) edge $(\mathcal{X}_{g},\mathcal{Y}_{g})$. Observe that $\mathcal{E}$ is a collection of disjoint cycles of length 2, or paths333 Assume that both $X$ and $Y$ are ordered according to $i$’s valuation: $v_{i}(x_{1})\geq\dots\geq v_{i}(x_{k})$ and the same for $Y$, where tie breaking is consistent among $X$ and $Y$. Observe that path between $\mathcal{Y}_{y_{j}}$ and $\mathcal{Y}_{y_{j^{\prime}}}$ implies that agent $i$ prefers $y_{j}$ over $y_{j^{\prime}}$, so a cycle of length greater than 2 is not possible. . Additionally, every vertex $v\in\mathcal{X}\cup\mathcal{Y}$ satisfies $deg(v)\in\{1,2\}$, where by $deg$ we consider both in- and out-degrees.
@@
Since $\mathcal{X}_{g^{\prime}}$ has no outgoing edges, $g^{\prime}\notin Y$. Also, since $g^{\prime}\in X$, then $g^{\prime}\notin G$. Thus, $g^{\prime}\in M\setminus(Y\cup G)$. It remains to show that $v_{i}(Y\cup\{g^{\prime}\})\geq\tau$. To this end, let $Q=\{g_{1},\dots,g_{\ell}\}\subseteq X\cap Y$, as defined by the path $P$. Observe that the set of goods which correspond to $\mathcal{Y}\cap P$ is exactly $Q\cup\{g\}$ (see Figure 1). Similarly, the set of goods corresponding to $\mathcal{X}\cap P$ is exactly $Q\cup\{g^{\prime}\}$. Since every vertex in $\mathcal{Y}\setminus P$ is pointed to by exactly one vertex in $\mathcal{X}\setminus P$, and $v_{i}(y_{j})\geq v_{i}(x_{j})$ for every $j\in[k]$, it follows that $v_{i}(Y\setminus(\{Q\cup\{g\}))\geq v_{i}(X\setminus(\{Q\cup\{g^{\prime}\}))$. By the last inequality and additivity of $v_{i}$ we get
@@
$\geq v_{i}(X\setminus(Q\cup\{g^{\prime}\}))+v_{i}(Q\cup\{g^{\prime}\})+v_{i}(\{g\})$
@@
which concludes the proof.
[…]
Figure 1: Illustration of the augmenting path argument in the proof of Lemma 7. The element $g\in G$ (whose corresponding node is $\mathcal{Y}_{g}$, $g=y_{1}$) is replaced with $g^{\prime}$ (whose corresponding node is $\mathcal{X}_{g^{\prime}}=\mathcal{X}_{4}$, $g^{\prime}=x_{4}$). We write $\mathcal{X}_{i}$ / $\mathcal{Y}_{i}$ rather than the cumbersome $\mathcal{X}_{x_{i}}$/$\mathcal{Y}_{y_{i}}$. It is purely an example that the path nodes are associated with a contiguous prefix of the goods, e.g., $\mathcal{X}_{1}$ could point to $\mathcal{Y}_{5}$.
Proof of Proposition 3.
@@
We will use a similar picking sequence procedure to the one given in Proposition 2, but with two rounds. First, we will only use the $m$ ‘original’ goods to determine the picking sequence, this is identical to the picking sequence presented earlier. Then, another picking sequence will use the ‘copies’ (and so will involve agents $1,\dots,t$, where each picks one good). Each agent in the second round may pick any good in $M$ which was not yet allocated in this round and was not picked by him in the first round. Let $A_{i}\subseteq M$ be the set of goods picked by agent $i$ in both rounds.
Any agent $i\in N\setminus\{1,\dots,t\}$ only participates in the first round, so as in Proposition 2, it holds that $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$. For agent $j\in\{1,\dots,t\}$, let $B_{j}$ be the set of $|A_{j}|-1$ goods he picked in the first round. To show that $v_{j}(A_{j})\geq v^{\prime}_{j}(A^{\prime}_{j})$, it is enough to show that at the beginning of the second round there is a set of goods $H\subseteq M\setminus B_{j}$, such that (i) $|H|\geq j$, (ii) each $h\in H$ satisfies $v_{j}(B_{j}\cup\{h\})\geq v^{\prime}_{j}(A^{\prime}_{j})$. This would imply that in round $j$, player $j$ can pick a good $h$ such that $A_{j}=B_{j}\cup\{h\}$ satisfies the claim.
Denote $A^{\prime}_{j}=\{j,j_{1},\dots,j_{k}\}$, $B_{j}=\{b_{1},\dots,b_{k}\}$, where $b_{\ell}$ was picked by agent $j$ in round $j_{\ell}$. The existence of $H$ follows from applying Lemma 7 with $X=\{\sigma_{j}(j_{1}),\dots,\sigma_{j}(j_{k})\}$, $Y=B_{j}$, $G=\{\sigma(g_{1}),\dots,\sigma_{j}(g_{j})\}$ and $\tau=v^{\prime}_{j}(A^{\prime}_{j})$. Observe that by definition of the first picking sequence, for any $\ell$, it holds that $v_{i}(b_{\ell})\geq v_{i}(\sigma_{j}(j_{\ell}))=v^{\prime}_{j}(j_{\ell})$. Also, $\tau=v^{\prime}_{j}(A^{\prime}_{j})=v^{\prime}_{j}(A^{\prime}_{j}\setminus\{g_{j}\})+v^{\prime}_{j}(g_{j})=v_{j}(X)+v_{j}(\sigma_{j}(g_{j}))=v_{j}(X\cup\{\sigma_{j}(g_{j})\})$. By definition, any good in $g\in G$ has $v_{j}(\{g\})\geq v_{j}(\sigma_{j}(g_{j}))$, and by additivity all the conditions of Lemma 7 are satisfied. This concludes the proof. ∎
For the rest of this section, we only consider ordered instances. Because the algorithms we use to solve these instances always output a simple allocation, and therefore, their guarantees, by Proposition 3, can be extended to any instance, even unordered.

### v3

@@ 4 Approximate Maximin Shares
It has been observed, for the original MMS problem (without copies), that it is without loss of generality to consider ordered instances [bouveret2016characterizing, barman2020approximation]. Notably, where copies are allowed, the same transformation as in [bouveret2016characterizing, barman2020approximation] may no longer yield a valid allocation, see example in Section C.1. To address this, we impose an additional constraint on the allocation in the ordered instance (see Definition 4.6), and show that under this constraint, it suffices to consider ordered instances.
Proposition 4.1.
[Reduction to ordered for simple allocations] Let $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ be a fair division instance and let $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ be its ordered counterpart. Let $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ be a simple allocation for $\mathcal{I}^{\prime}$ with $t$ distinct copies. Then, $A^{\prime}$ can be converted in poly-time to an allocation for $\mathcal{I}$ with $t$ distinct copies, $A=(A_{1},\dots,A_{n})$, such that for every agent $i$, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$.
We then devise algorithms that produce allocations adhering to this constraint.
In Section 4.2, we review some of the known reduction rules and introduce new ones. Then, in Section 4.3, we present a simple algorithm $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ which combines the idea of bag-filling and round robin to achieve an $\alpha$-MMS allocation with copies (see Algorithm 4). This algorithm is the building block for achieving $6/7$-MMS and $4/5$-MMS with copies in the sections that follow.
Theorem 4.2.
Given an instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ with $n^{*}$ agents and additive valuations, for $\alpha\leq 6/7$, there exists an $\alpha$-MMS allocation with at most $\lfloor n^{*}/2\rfloor$ many distinct copies.
Theorem 4.3.
Given an instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ with $n^{*}>5$ agents and additive valuations, for $\alpha\leq 4/5$, there exists an $\alpha$-MMS allocation with at most $\lfloor n^{*}/3\rfloor$ many distinct copies.
The missing proofs of this section can be found in Appendix C.
@@ 4.1 Ordered Instance
Given an instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, for every agent $i\in N$, let $\sigma_{i}:[m]\to[m]$ be a permutation that orders the goods according to $i$’s preferences, namely $\sigma_{i}(j)$ is the $j$th most valued good (breaking ties by index [^2] Any consistent tie breaking rule will do., and thus $v_{i}(\sigma_{i}(1))\geq v_{i}(\sigma_{i}(2))\geq\dots\geq v_{i}(\sigma_{i}(m))$.
As done by Bouveret and Lemaître [bouveret2016characterizing], we construct an ordered instance $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ with a new set of “virtual goods”, $M^{\prime}=\{1,\dots,m\}$, and modified valuations, $\{v^{\prime}_{i}\}_{i\in N}$, such that $v^{\prime}_{i}(j)=v_{i}(\sigma_{i}(j))$ for any $j\in M^{\prime}$.
Definition 4.4 (Ordered Counterpart ([bouveret2016characterizing])).
Given a fair division instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, its ordered counterpart, $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$, is an instance with the same set of agents and a set of “virtual goods”, $M^{\prime}$, of the same cardinality as $M$. For each agent $i\in[n]$, and any $j\in M^{\prime}$, it holds that $v^{\prime}_{i}(j)=v_{i}(\sigma_{i}(j))$.
Observe that by definition, all agent in $\mathcal{I}^{\prime}$ have the same ordinal preferences, i.e., for every $i\in N$, $v^{\prime}_{i}(1)\geq v^{\prime}_{i}(2)\geq\dots\geq v^{\prime}_{i}(m)$. It was shown in [bouveret2016characterizing] that to guarantee agent $i$ a value of at least $\alpha_{i}$ in an instance $\mathcal{I}$, it suffices to find an allocation providing this guarantee in its ordered counterpart $\mathcal{I}^{\prime}$. This is cast in the following proposition.
Proposition 4.5 (Reduction to ordered without copies [bouveret2016characterizing]).
Let $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ be a fair division instance and let $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ be its ordered counterpart. Any allocation (without copies) $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ for $\mathcal{I}^{\prime}$ can be converted in poly-time to an allocation (without copies) $A=(A_{1},\dots,A_{n})$ for $\mathcal{I}$, such that for every agent $i$, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$.
@@
Since by iteration $j$ exactly j1j-1 items have been chosen, agent $i$ always chooses a good $g\in M$ such that $v_{i}(g)\geq v_{i}(\sigma_{i}(j))=v^{\prime}_{i}(j)$ and thus, by additivity, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$. ∎
We will show a similar result to Proposition 4.5, for allocations with $t$ distinct copies in which the top $t$ goods (and their copies) are allocated to $2t$ different agents.
Definition 4.6 (Simple Allocation).
@@
  • There is at most one duplication of each good.
@@
Recall that agent $i$ cannot get two instances of the same good in $A^{\prime}_{i}$. [^3] Otherwise, and if additive valuations are extended so that two instances of the same good yield twice its value, then the same argument as in Proposition 4.5 would have implied the correctness of Proposition 4.1.
See 4.1
In order to prove Proposition 4.1, we presented the following useful lemma.
Lemma 4.7.
@@
If there exists a set $G\subset M\setminus X$ of $r$ distinct goods such that $v_{i}(X\cup\{g\})\geq\tau$ for all $g\in G$, then there exists a set $H\subset M\setminus Y$ of $r$ distinct goods such that $v_{i}(Y\cup\{h\})\geq\tau$ for all $h\in H$.
@@
For every $j\in[k]$, we add the (black) edge $(\mathcal{Y}_{j},\mathcal{X}_{j})$. In addition, for any good $g\in X\cap Y$, we add the (red) edge $(\mathcal{X}_{g},\mathcal{Y}_{g})$. Observe that $\mathcal{E}$ is a collection of disjoint cycles of length 2, and paths [^4] Assume that both $X$ and $Y$ are ordered according to $i$’s valuation: $v_{i}(x_{1})\geq\dots\geq v_{i}(x_{k})$ and the same for $Y$, where tie breaking is consistent among $X$ and $Y$. Observe that path between $\mathcal{Y}_{y_{j}}$ and $\mathcal{Y}_{y_{j^{\prime}}}$ implies that agent $i$ prefers $y_{j}$ over $y_{j^{\prime}}$, so a cycle of length greater than 2 is not possible. . Additionally, every vertex $v\in\mathcal{X}\cup\mathcal{Y}$ satisfies $deg(v)\in\{1,2\}$, where by degdeg we consider both in- and out-degrees.
@@
Since $\mathcal{X}_{g^{\prime}}$ has no outgoing edges, $g^{\prime}\notin Y$. Also, since $g^{\prime}\in X$, then $g^{\prime}\notin G$. Thus, $g^{\prime}\in M\setminus(Y\cup G)$. It remains to show that $v_{i}(Y\cup\{g^{\prime}\})\geq\tau$. To this end, let $Q=\{g_{1},\dots,g_{\ell}\}\subseteq X\cap Y$, as defined by the path $P$. Observe that the set of goods which correspond to $\mathcal{Y}\cap P$ is exactly $Q\cup\{g\}$ (see Figure 1). Similarly, the set of goods corresponding to $\mathcal{X}\cap P$ is exactly $Q\cup\{g^{\prime}\}$. Observe that the vertices in $P$ can be thought of as pairs where each vertex in $\mathcal{Y}\cap P$ has value $\geq$ than its succeeding vertex in $\mathcal{X}\cap P$. Thus, when removing these edges from the graph the property $v_{i}(y_{j})\geq v_{i}(x_{j})$ is maintained and thus $v_{i}(Y\setminus(\{Q\cup\{g\}))\geq v_{i}(X\setminus(\{Q\cup\{g^{\prime}\}))$. By the last inequality and additivity of $v_{i}$ we get
@@
$\geq v_{i}(X\setminus(Q\cup\{g^{\prime}\}))+v_{i}(QWe\cup\{g^{\prime}\})+v_{i}(\{g\})$
@@
which concludes the proof. ∎
$\mathcal{X}_{{1}}$$\mathcal{X}_{{2}}$$\mathcal{X}_{{3}}$$\mathcal{X}_{{4}}$$\mathcal{X}_{{5}}$$\mathcal{Y}_{{1}}$$\mathcal{Y}_{{2}}$$\mathcal{Y}_{{3}}$$\mathcal{Y}_{{4}}$$\mathcal{Y}_{{5}}\cdots\cdotsgggg^{\prime}$QQQQ
Figure 1: Illustration of the augmenting path argument in the proof of Lemma 4.7. The element $g\in G$ (whose corresponding node is $\mathcal{Y}_{g}$, $g=y_{1}$) is replaced with $g^{\prime}$ (whose corresponding node is $\mathcal{X}_{g^{\prime}}=\mathcal{X}_{4}$, $g^{\prime}=x_{4}$). We write $\mathcal{X}_{i}$ / $\mathcal{Y}_{i}$ rather than the cumbersome $\mathcal{X}_{x_{i}}$$\mathcal{Y}_{y_{i}}$. It is purely an example that the path nodes are associated with a contiguous prefix of the goods, e.g., $\mathcal{X}_{1}$ could point to $\mathcal{Y}_{5}$.
Proof of Proposition 4.1.
@@
We will use a similar picking sequence procedure to the one given in Proposition 4.5, but with two rounds. First, we will only use the $m$ ‘original’ goods to determine the picking sequence, this is identical to the picking sequence presented earlier. Then, another picking sequence will use the ‘duplications’ (and so will involve agents $1,\dots,t$, where each picks one good). Each agent in the second round may pick any good in $M$ which was not yet allocated in this round and was not picked by him in the first round. Let $A_{i}\subseteq M$ be the set of goods picked by agent $i$ in both rounds.
Any agent $i\in N\setminus\{1,\dots,t\}$ only participates in the first round, so as in Proposition 4.5, it holds that $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$. For agent $j\in\{1,\dots,t\}$, let $B_{j}$ be the set of $|A_{j}|-1$ goods he picked in the first round. To show that $v_{j}(A_{j})\geq v^{\prime}_{j}(A^{\prime}_{j})$, it is enough to show that at the beginning of the second round there is a set of goods $H\subseteq M\setminus B_{j}$, such that (i) $|H|\geq j$, (ii) each $h\in H$ satisfies $v_{j}(B_{j}\cup\{h\})\geq v^{\prime}_{j}(A^{\prime}_{j})$. This would imply that in round $j$, player $j$ can pick a good $h$ such that $A_{j}=B_{j}\cup\{h\}$ satisfies the claim.
Denote $A^{\prime}_{j}=\{j,j_{1},\dots,j_{k}\}$, $B_{j}=\{b_{1},\dots,b_{k}\}$, where $b_{\ell}$ was picked by agent $j$ in round $j_{\ell}$. The existence of $H$ follows from applying Lemma 4.7 with $X=\{\sigma_{j}(j_{1}),\dots,\sigma_{j}(j_{k})\}$, $Y=B_{j}$, $G=\{\sigma(g_{1}),\dots,\sigma_{j}(g_{j})\}$ and $\tau=v^{\prime}_{j}(A^{\prime}_{j})$. Observe that by definition of the first picking sequence, for any $\ell$, it holds that $v_{i}(b_{\ell})\geq v_{i}(\sigma_{j}(j_{\ell}))=v^{\prime}_{j}(j_{\ell})$. Also, $\tau=v^{\prime}_{j}(A^{\prime}_{j})=v^{\prime}_{j}(A^{\prime}_{j}\setminus\{g_{j}\})+v^{\prime}_{j}(g_{j})=v_{j}(X)+v_{j}(\sigma_{j}(g_{j}))=v_{j}(X\cup\{\sigma_{j}(g_{j})\})$. By definition, any good in $g\in G$ has $v_{j}(\{g\})\geq v_{j}(\sigma_{j}(g_{j}))$, and by additivity all the conditions of Lemma 4.7 are satisfied. This concludes the proof. ∎
For the rest of this section, we only consider ordered instances. Because the algorithms we use to solve these instances always output a simple allocation, their guarantees, by Proposition 4.1, can be extended to any instance, even unordered.